• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Carlos

    Log entry time 06:29:28 on October 7,2010

    Entry number 333873

    keyword=Owl shift summary

    Shift started with accelerator down due to a power supply that died.

    Beam got back at around 00:40 and Compton tuning was resumed.

    Around 4:00 Compton tunning was stopped, and we started taking some beam on target with the dipole at 200A.

    We checked the spot on the BeO, and moved to carbon-hole with raster. We saw the hole with spot++.

    We took one "commissioning" run (7048) with 5uA on 15cm LH2, with T3 as a trigger. We saw DC signals from the calorimeter anode currents. This run can also be used to test lumis. I'm replaying this run right now to check HRS detectors.

    At 6:00, we requested to go into controlled access.